OpenStudy (saifoo.khan):

What are the types of method to solve substitution equations with 3 variables?

OpenStudy (saifoo.khan):

I just want the Names.

OpenStudy (saifoo.khan):

Or it's simply known as "substution method"?

OpenStudy (saifoo.khan):

substitution*

OpenStudy (agreene):

trigonometric substitution basic substitution complex substitution ??

OpenStudy (turingtest):

Gaussian-Elimination, Cramer's Rule, L-U breakdown. There are countless ways to do these. Take a Linear Algebra course if you want to know more.

OpenStudy (saifoo.khan):

o_O

OpenStudy (saifoo.khan):

x – 2y + z = 0 2x – 3y – 4z = –9 x + 2y – 5z = 0 Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point)

OpenStudy (saifoo.khan):

i solve them by eliminating z.. and so on..

OpenStudy (agreene):

I would use method of elimination. Thats what thats called hehe

OpenStudy (saifoo.khan):

what about part 2? "WHY" would u use it? LOL

OpenStudy (turingtest):

you can see that adding eqn.1 to eqn.3 the y's will disappear. and you'll have 2x-4z=0 which will help you in the next step.

OpenStudy (jamesj):

I'd use Gaussian elimination because the method is elegant, very transparent and avoids the confusion that sometimes arise when using substitution (which equation have I substituted into already?; what equation should I use next?).